### v1

# Active nematic response to a deformable body or boundary: elastic deformations and anchoring-induced flow

Abstract
A body immersed in a nematic liquid crystal disturbs the fluid’s preferred molecular configuration and increases its stored elastic energy. In an active nematic, the fluid components also generate a stress in the bulk fluid. By introducing either an immersed body or boundary, a fluid flow can be triggered due to anchoring boundary conditions. The fluid imposes elastic, viscous, and active stresses on such surfaces, which, if compliant, may result in a surface deformation. We study the deformations and stress in a linearly elastic body placed in an active nematic in two-dimensions. Using complex variables techniques, exact expressions for the fluid flow, director field, surface tractions, and body deformation are derived. Qualitative differences between elastic and active stress-driven deformations are identified, depending on an active Ericksen number, anchoring conditions, and body material properties, thereby suggesting a new method for measuring the active stresses in active anisotropic biological systems. Flow profiles, external confinement, and anchoring-induced stirring are also addressed.
The eukaryotic cell is host to a wide variety of deformable organelles, from the nucleus to mitochondria [1, 2, 3, 4, 5, 6]. These organelles exist in a complex, potentially anisotropic, viscous environment [7, 8], and experience mechanical stresses due to actively driven cytoplasmic flows [9]. The cell membrane itself is also deformable, and is commonly remodelled by internal processes. For instance, active anisotropic stresses are of great functional importance in cytoskeletal remodeling of the nucleus and chromatin configurations [10, 11], and in the metaphase spindle [12, 13, 14]. On a larger length scale, active systems like bacterial swarms and biofilms also exhibit anisotropy, among other complex rheological properties [15, 16, 17, 18]. Understanding how such bodies affect, and are affected by, this complex environment is of generic importance.
Even when free of bodies, actively-driven anisotropic flows exhibit very rich and complex dynamics [19, 20, 21, 22, 23]. Stability [24, 21, 25, 26, 27, 28, 29, 30], topology [31], pattern formation [32, 33, 34], and mixing [35, 36] have seen substantial theoretical treatment. Confinement introduces an additional length scale, which can affect the emergent spatial structure [37, 38]. Strong active nematic stresses can also change the shape of deformable confining surfaces [39, 40, 41, 42, 43, 44, 45, 46], and they play a prominent role in directing the motility of cells [47, 48] and active droplets [49, 50, 51, 52].
When an active nematic contains an immersed body, or inclusion, additional elastic stresses are immediately introduced. In passive systems, these stresses result in intricate and beautiful textures [53, 54, 55], and can be used to measure cell material properties [56, 57]. In the presence of activity, the combination of fluid elasticity and anisotropy redirects actively driven flows, leading to complex immersed body dynamics. Recent examples include an investigation of the biased rotation of gear-like bodies [58, 59], and transitions from fixed point to limit cycle to chaotic dynamics of a circular disk [60]. How body deformability affects, and is affected, by such an active environment has not been well explored.
In this letter, we present a model for an active anisotropic fluid with a soft internal or external boundary, confined to two-dimensions (2D). The model integrates Ericksen–Leslie theory [61, 62, 63] for active nematic liquid crystals (LCs), a common model for active anisotropic fluid [64, 23], and linear elasticity for deformable solids. Complex variable techniques yield novel analytical representations for the director field, velocity field, and boundary deformation. We demonstrate this for both an immersed and bounding cylinder, highlighting the insights that these analytical solutions provide.
Active nematic LCs can be described by their locally-averaged molecular orientation, $\bm{n}(\bm{x},t)$, and fluid velocity $\bm{u}(\bm{x},t)$, with spatial position $\bm{x}$, time $t$, and $|\bm{n}|=1$. The director field, $\bm{n}$, evolves due to gradients in the velocity field, with strain rate tensor $\bm{\mathrm{E}}=(\nabla\bm{u}+\nabla\bm{u}^{T})/2$ and vorticity tensor $\bm{\Omega}=(\nabla\bm{u}-\nabla\bm{u}^{T})/2$, and a relaxation towards equilibrium under the action of the molecular field, $\bm{h}$. In the one-constant approximation, $\bm{h}=K\nabla^{2}\bm{n}$, for the single (Frank) elastic constant $K$. Altogether,
$\partial_{t}\bm{n}+\bm{u}\cdot\nabla\bm{n}=\bm{n}\cdot\bm{\Omega}+\lambda\bm{n}\cdot\bm{\mathrm{E}}\cdot(\bm{\mathrm{I}}-\bm{n}\bm{n})+\tilde{\bm{h}}/\gamma,$ (1)
with $\gamma$ rotational viscosity, $\lambda$ tumbling/reactive parameter, and $\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\bm{h}$ the transverse part of the molecular field, where $\bm{n}\bm{n}$ is a dyadic product [61, 62, 63]. Assuming incompressibility, the fluid velocity satisfies the momentum and mass conservation equations
$\nabla\cdot\left(\bm{\sigma}^{e}+\bm{\sigma}^{v}+\bm{\sigma}^{a}-p\bm{\mathrm{I}}\right)=\bm{0}\quad\text{and}\quad\nabla\cdot\bm{u}=0,$ (2a)
for the pressure $p$ and elastic, viscous, and active stress tensors
$\bm{\sigma}^{e}$ $=-K\nabla\bm{n}\cdot\nabla\bm{n}^{T}+\frac{1}{2}(1-\lambda)\bm{n}\bm{\tilde{h}}-\frac{1}{2}(1+\lambda)\bm{\tilde{h}}\bm{n},$ (3a)
[…]
$\bm{\sigma}^{a}$ $=2\alpha\bm{n}\bm{n},$ (3c)
respectively. Here, $\mu$ is the solvent viscosity, $\mu_{1}$ and $\mu_{2}$ are anistropic viscosities, and $\alpha$ is the activity strength, which can describe both extensile ($\alpha<0$) and contractile ($\alpha>0$) activity [64, 65, 23].
We non-dimensionalize the system by using an intrinsic length scale $L$ (e.g. the size of an immmersed body), velocity scale $U$ (to be defined), and stress scale $K/L^{2}$. The dimensionless system is then described by the dimensionless Ericksen number, activity strength, and viscosity ratios [66]:
$\mathrm{Er}\coloneqq\frac{\mu UL}{K},\quad\mathit{A}\coloneqq\frac{\alpha L}{\mu U},\quad\gamma^{\prime}\coloneqq\frac{\gamma}{\mu},\quad\mu_{j}^{\prime}\coloneqq\frac{\mu_{j}}{\mu}.$ (4a)
For small rotational Ericksen numbers, $\gamma^{\prime}\mathrm{Er}\ll 1$, the fluid flow does not affect the director field and the molecular field relaxes to equilibrium. The director evolution equation (1) takes the dimensionless form
$\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\nabla^{2}\bm{n}=\mathcal{O}(\gamma^{\prime}\mathrm{Er}).$ (5)
By further assuming small viscosity ratios, $\gamma^{\prime},\mu_{i}^{\prime}\ll 1$, the momentum equation (2a) takes the dimensionless form
$\nabla^{2}\bm{u}-\nabla p=\frac{1}{\mathrm{Er}}\nabla\cdot\bigl{(}\nabla\bm{n}\cdot\nabla\bm{n}^{T}\bigr{)}-2\mathit{A}\nabla\cdot(\bm{n}\bm{n}),$ (6)
up to $\mathcal{O}(\gamma^{\prime},\mu_{i}^{\prime})$. This describes an isotropic Stokes flow with active and elastic driving forces [66].
In the examples presented in this letter, the fluid flow is induced by activity alone; thus, the velocity scale is set by the activity strength, $U=|\alpha|L/\mu$, and the assumption of a small Ericksen number is equivalent to assuming weak activity, $\gamma^{\prime}\mathrm{Er}=|\alpha|\gamma L^{2}/(\mu K)\ll 1$. Furthermore, in 2D, the assumption of a small viscosity ratio $\mu_{2}^{\prime}$ is unnecessary since it only corresponds to an anisotropy out-of-the-plane. Instead, $\mu_{2}$ can be absorbed into the isotropic viscosity by taking $\mu\mapsto\mu-\mu_{2}/2$.
Figure 1: The activity-induced flow of a 2D nematic LC outside a cylinder for $\mathit{A}=-1$ (i.e. extensile activity). (a) The director lines for $w=10$ are shown as blue curves. (b,c) The activity–induced velocity field is shown for (b) $w=1$ and (c) $w=10$. The black arrows show the flow direction and the color denotes the fluid speed, $\sqrt{u^{2}+v^{2}}$. Contractile activity with $\mathit{A}=1$ corresponds to flipping the flow direction.
In two dimensions, the director field can be described by a director angle $\theta(x,y)\in[0,\pi)$ such that $\bm{n}=(\cos\theta,\sin\theta)$. Additionally, the conservation of mass (2b) implies that a streamfunction $\psi(x,y)$ can be introduced such that $\bm{u}=(\partial_{y}\psi,-\partial_{x}\psi)$. In this case, the equilibrium equation (5) reduces to Laplace’s equation, $\nabla^{2}\theta=0$, while the momentum equation (6) is equivalent to a Poisson’s equation for the scalar vorticity, $\omega=-\nabla^{2}\psi$. Here, we absorb the isotropic parts of the stress tensors into the pressure by taking $p\mapsto p+\mathit{A}-|\nabla\theta|^{2}/(2\mathrm{Er})$.
At any boundaries, the fluid velocity is assumed to vanish, while the director relaxes to some preferential angle $\phi$ with some anchoring strength $W$ (i.e. Rapini–Papoular anchoring [67]):
$\partial_{x}\psi=\partial_{y}\psi=0\hskip 6.00006pt\text{and}\hskip 6.00006pt\partial_{\nu}\theta=(w/2)\sin\left[2(\theta-\phi)\right],$ (7a)
for the normal derivative pointing into the LC $\partial_{\nu}=\bm{\hat{\nu}}\cdot\nabla$ and the dimensionless anchoring strength $w\coloneqq WL/K.$
Consider the complex position $z=x+\mathrm{i}y$ and its complex conjugate $\bar{z}=x-\mathrm{i}y$. Using the Wirtinger derivatives, $\partial_{x}=\partial_{z}+\partial_{\bar{z}}$ and $\partial_{y}=\mathrm{i}\partial_{z}-\mathrm{i}\partial_{\bar{z}}$, one can make a change of variables from $(x,y)$ to $(z,\bar{z})$ [68]. In these complex coordinates, the system takes the form
$\partial_{z\bar{z}}\theta=0\quad\text{and}\quad\partial_{\bar{z}}(p-\mathrm{i}\omega)=\mathit{A}\partial_{z}e^{2\mathrm{i}\theta},$ (8a)
with vorticity $\omega=-4\partial_{z\bar{z}}\psi$ and velocity, $\bm{u}=(u,v)$, given by $u-\mathrm{i}v=2\mathrm{i}\partial_{z}\psi$. Integrating (8) yields
$\theta$ $=-\arg f^{\prime}(z)$ (9a)
$\text{and}\quad p-\mathrm{i}\omega$ $=\mathit{A}\partial_{z}\bigl{[}g(z)+\overline{f(z)}/f^{\prime}(z)\bigr{]},$ (9b)
for some locally-holomorphic functions $f$ and $g$ [68]. (Throughout, we shall assume the LC does not contain any topological defects, and so, $f$ is non-zero and singularity free inside the LC.) Inserting the imaginary part of (9b) into $\partial_{z\bar{z}}\psi=-\omega/4$ and integrating yields
$\psi=\frac{\mathit{A}}{4}\operatorname{Im}\Bigl{[}\bar{z}g(z)+h(z)+\overline{\int f(z)\mathop{}\!\mathrm{d}z}\,/f^{\prime}(z)\Bigr{]},$ (10)
for another locally-holomorphic function $h$. The first two terms in (10) correspond to the well-known form of a biharmonic function with Goursat functions $h$ and $g$ [69]. The last term is a particular solution that accounts for the active forcing.
Analytical solutions for the director and velocity fields are thus obtained, provided that functions $f$, $g$, and $h$, which are locally holomorphic within the LC and satisfy the boundary conditions in (7), may be found. The problem for $f$ derived here is equivalent to the potential problem explored in Refs. [70, 71] for a passive LC. It is important to note that the solutions (9)–(10) are not invariant under a conformal mapping since they contain a mixture of holomorphic (function of $z$) and antiholomorphic (function of $\bar{z}$) terms.
As an example, consider a cylinder of radius $L$ of infinite extent immersed in a weakly-active nematic LC. In the far-field, we assume the LC is aligned horizontally in a direction orthogonal to the cylinder’s long axis and the fluid velocity vanishes. Additionally, the LC is assumed to be tangentially anchored to the cylinder with a finite strength, i.e. (7b) with $\phi=\arg(\mathrm{i}z)\mod\pi$ on $|z|=1$. Without the body, the system sits at an equilibrium, with $\bm{n}=\bm{\hat{x}}$ and $\bm{u}=\bm{0}$, and the active stress is absorbed by the pressure.
First, we determine the LC director angle $\theta=-\arg f^{\prime}$. Here, the potential $f(z)$ must be locally holomorphic in $1<|z|<\infty$ and satisfy the finite anchoring condition on $|z|=1$, and $f(z)\sim z$ as $|z|\to\infty$. The solution to this problem that minimizes the Frank elastic energy is, for any $w>0$,
$\theta=-\arg f^{\prime}(z)=-\arg(1-\rho^{2}/z^{2}),$ (11)
where $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$ is an effective radius [70]. The corresponding director lines (which lay parallel to $\bm{n}$) are plotted in Fig. 1(a) for $w=10$.
Next, with (11), we determine the activity-induced fluid flow. Here, the Goursat functions $h(z)$ and $g(z)$ must be locally holomorphic in $1<|z|<\infty$ and yield a fluid velocity that vanishes on $|z|=1$ and as $|z|\to\infty$, ultimately yielding the streamfunction
$\psi=\frac{\mathit{A}\rho^{2}}{8}\operatorname{Im}\biggl{[}\frac{(|z|^{2}-1)^{2}}{z^{2}(z^{2}-\rho^{2})}+\frac{1-|z|^{2}+2\log|z|}{(z^{2}-\rho^{2})/(2\rho^{2})}\biggr{]},$ (12)
see [66]. The corresponding velocity field is shown in Fig. 1(b,c) for $w=1$ and $w=10$, with $\mathit{A}=-1$ (i.e. extensile activity). A prominent feature is a plume extending symmetrically to both the left and right of the cylinder. Far from the body, the velocity field behaves as
$u+\mathrm{i}v\sim\bigl{[}\rho^{2}\cos(2\alpha)-\cos(4\alpha)\bigr{]}\mathit{A}\rho^{2}e^{\mathrm{i}\alpha}/(2|z|),$ (13)
as $|z|\to\infty$, for $\alpha=\arg z$; the plume angle thus decreases from $\pi/4$ to 00 as the anchoring strength $w$ is increased from 00 to $\infty$, while the plume speed is largest at $w=8/3$. The velocity field does not appear as a dipole or stresslet in the far-field — the mere presence of the body initiates additional flow far away due to the distributed active stress, leading to bulk flow release. This also demonstrates that a body with non-trivial anchoring conditions is sufficient to induce a large scale flow through its presence alone.
Both the elasticity associated with passive LCs and the flow induced by activity generate stresses that act on any boundaries. Here, the passive LC introduces a surface traction $\bm{t}^{e}=\bm{\hat{\nu}}\cdot\hat{\bm{\sigma}}^{e}+\partial_{s}\bm{t}^{s}$, where $\hat{\bm{\sigma}}^{e}=(1/2)\lvert\nabla\theta\rvert^{2}\bm{\mathrm{I}}-\nabla\theta\nabla\theta$ is the (traceless) elastic stress tensor and $\bm{t}^{s}$ is a surface stress vector caused by the finite surface-anchoring (7b) [66]. The activity generates an additional surface traction $\bm{t}^{a}=\mathrm{Er}\,\bm{\hat{\nu}}\cdot(\hat{\bm{\sigma}}^{v}+\mathit{A}\hat{\bm{\sigma}}^{a}-p\bm{\mathrm{I}})$ for the (traceless) viscous and active stress tensors $\hat{\bm{\sigma}}^{v}=\nabla\bm{u}+\nabla\bm{u}^{T}$ and $\hat{\bm{\sigma}}^{a}=2\bm{n}\bm{n}-\bm{\mathrm{I}}$, respectively.
If a boundary is soft, it will generally be deformed by the surface tractions above. We assume the boundary is part of an isotropic solid that deforms according to plane stress/strain in the framework of linear elasticity, with shear modulus $\mu_{s}$ and Poisson’s ratio $\nu$ [72, 73]. In this framework, the deformed (Eulerian) frame and the reference (Lagrangian) frame are mathematically equivalent. At equilbrium, the elastic solid can be described by an Airy-stress function, $\mathcal{A}(x,y)$, which satisfies the biharmonic equation, $\nabla^{4}\mathcal{A}=0$. We can, thus, write
$\mathcal{A}=\operatorname{Im}[\bar{z}G(z)+H(z)],$ (14)
for two (new) Goursat functions, $G$ and $H$, that are locally holomorphic within the solid (but not necessarily the LC), analogous to $g$ and $h$ in (10) [72]. These Goursat functions yield the stress tensor within the solid, $\bm{\Sigma}$, via the Kolosov–Muskhelishvili formulae [66]. Additionally, the solid displacement, $(U,V)$, can be expressed as
$2M(U-\mathrm{i}V)=\mathrm{i}\kappa\overline{G(z)}+\mathrm{i}\bar{z}G^{\prime}(z)+\mathrm{i}H^{\prime}(z),$ (15)
up to an arbitrary body displacement and rotation, for the dimensionless elastic modulus $\mathit{M}\coloneqq\mu_{s}L^{2}/K$ and $\kappa=(3-4\nu)$ or $(3-\nu)/(1+\nu)$ for plane strain or plane stress, respectively.
The LC–solid interface must also be at equilibrium, and so, the stresses acting on it must balance: $-\bm{\hat{\nu}}\cdot\bm{\Sigma}=\bm{t}^{e}+\bm{t}^{a}$. Provided we can find functions $G$ and $H$ that are locally holomorphic within the solid and satisfy this stress balance, we will have obtained analytic solutions for the solid’s deformation. However, yet again, the above solution is not invariant under a conformal map, and analytical progress can only be made in specific cases.
We return to the example of a unit cylinder subject to (finite) tangential anchoring. Here, the director angle (11) and streamfunction (12) yield analytic expressions for the elastic traction, $t^{e}_{x}+\mathrm{i}t_{y}^{e}$, and activity-induced traction, $t^{a}_{x}+\mathrm{i}t_{y}^{a}=\mathit{A}\mathrm{Er}(1/z-\rho^{2}z)$, acting on the cylinder. The cylinder is now assumed to be deformable, and so, we seek the Airy stress (14). The Goursat functions $G(z)$ and $H(z)$ must be locally holomorphic in $|z|<1$ and satisfy the stress balance condition on $|z|=1$. See [66] for the solution to this problem.
For passive LCs ($\mathit{A}=0$), the deformation acts to reduce the elastic energy stored in the LC by flattening the cylinder. Figure  2 shows the LC configurations and body deformations for three different anchoring strengths and $\kappa=5/3$ (i.e. an incompressible cylinder, $\nu=1/2$, under plane stress). For large anchoring strengths, $w\gg 1$, most of the LC energy is stored local to the left and right poles of the cylinder, with a singular energy emerging in the case of strong (infinite) anchoring ($w=\infty$) due to the appearance of two $-1$ Boojum topological defects at $z=\pm 1$ [74, 70]. This localization of energy within the LC results in large localized surface tractions: $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(w^{2})$ for $z\pm 1=\mathcal{O}(1/w)$, and $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(1)$ otherwise, as $w\to\infty$. This, in turn, leads to large localized deformations of the cylinder, see Fig. 2. In particular, we find that
$U+\mathrm{i}V\sim\frac{3w}{8M}\bigl{[}\mathcal{U}(1-z)-\mathcal{U}(1+z)\bigr{]},$ (16)
as $w\to\infty$, where
$\mathcal{U}(Z/w)=\kappa/(1+Z)+1/(1+\bar{Z})+(Z+\bar{Z})/(1+\bar{Z})^{2},$ (17)
corresponds to a deformation that is only apparent local to the left and right poles, i.e. when $z\pm 1=\mathcal{O}(1/w)$. As the anchoring strength decreases, stress is no longer concentrated at the poles and the maximum shear stress $\tau(x,y)=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$ (a common indicator of mechanical failure [73]) is greatest at (±$(\pm x_{\max},0)$, where $x_{\max}$ decreases from $1$ for $w=\infty$, to 00 for $w\leq 24/35$, see Fig. 2.
Figure 2: (top) Deformation of soft cylinders immersed in passive nematic LCs with $\mathit{A}\mathrm{Er}=0$, $\mathit{M}/w=5$, $\kappa=5/3$, and dimensionless anchoring strengths ① $w=0.1$, ② $w=1$, and ③ $w=10$. Color indicates the maximum shear stress, $\tau=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The positions, $(x_{\max},0)$, and value, $\tau_{\max}$, of the largest maximum shear stress plotted as a function of $w$.
For active LCs ($A\neq 0$), the activity leads to an additional deformation:
$U+\mathrm{i}V=U^{e}+\mathrm{i}V^{e}+\mathit{A}\mathrm{Er}\bigl{[}(1-\kappa)\rho^{2}z+2\bar{z}\bigr{]}/(4\mathit{M}),$ (18)
where $(U^{e},V^{e})$ is the deformation of a cylinder immersed in a passive LC. As this addition is proportional to the activity strength, $\mathit{A}$, activity can act with or against the elastic effects of the LC, depending on if it is extensile ($\mathit{A}<0$) or contractile ($\mathit{A}>0$). In particular, the aspect ratio, $\mathcal{A}_{R}$, of the deformed cylinder is given by
$\mathcal{A}_{R}\sim 1+U\bigr{|}_{z=1}-V\bigr{|}_{z=\mathrm{i}}=1+(\mathit{A}\mathrm{Er}-\mathcal{W})/\mathit{M},$ (19)
where
$\mathcal{W}=\frac{3\kappa}{2\rho}\left(\operatorname{arctanh}\rho-\arctan\rho\right)-\frac{3w}{8}(1+\kappa).$ (20)
Thus, if $\mathit{A}\mathrm{Er}>\mathcal{W}$ (which includes passive LCs since $\mathcal{W}<0$), the deformed shape is elongated with the $x$-axis (i.e. the preferred axis of the LC). While, if $\mathit{A}\mathrm{Er}<\mathcal{W}$, the shape is elongated with the $y$-axis (i.e. perpendicular to the preferred axis). The aspect ratio is one at the critical value $\mathit{A}\mathrm{Er}=\mathcal{W}$, see Fig. 3. Therefore, contractile activity ($\mathit{A}>0$), in this setting, always leads to parallel elongation, while extensile activity ($\mathit{A}<0$) yields perpendicular elongation, as long as the activity is sufficiently strong.
Figure 3: (top) Deformation of soft cylinders immersed in active nematic LCs with $w=5$, $\mathit{M}=50$, $\kappa=5/3$, and ④ $\mathit{A}\mathrm{Er}=-8<\mathcal{W}$, ⑤ $\mathit{A}\mathrm{Er}=\mathcal{W}\approx-3.55$, and ⑥ $\mathit{A}\mathrm{Er}=4>\mathcal{W}$. Color indicates the maximal shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines. (bottom) The critical active Ericksen number, $\mathit{A}\mathrm{Er}=\mathcal{W}$, at which the deformed cylinder has aspect ratio one, plotted as a function of $w$ for $\kappa=5/3$. The points ①–③ correspond to the configurations shown in Fig. 2.
The fluid flow triggered by the presence of a boundary may also be advantageous in biological cells, settings in which a disperse distribution of molecules, such as nutrients, chemicals, or oxygen, benefit the system’s function [75, 76, 77]. Consider an active nematic LC bounded inside, and tangentially anchored to, a cylinder with anchoring strength $w$. Here, the director angle of the LC can be written as
$\theta=\arg(1-\rho^{2}z^{2}),$ (21)
for the effective radius introduced in the external problem, $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$, while the activity-induced flow can be expressed in terms of a streamfunction
$\begin{split}\psi=\frac{\mathit{A}}{8}(|z|^{2}-1)\operatorname{Im}\Big{[}&\frac{1-\rho z}{z^{2}}\log(1-\rho z)\\ +&\frac{1+\rho z}{z^{2}}\log(1+\rho z)\Big{]},\end{split}$ (22)
see [66]. As before, the elasticity, activity, and viscosity of the LC will induce surface stresses on the boundary, which, if compliant, will deform. Figure 4(a) shows the director lines inside a cylinder that has deformed according to linear elasticity [66]. The final deformed shape resembles a tactoid [78], with a localized deformation at the left and right poles occurring for large anchoring strengths due to the appearance of two $+1$ aster defects when $w=\infty$. The deformed shape is dependent on the active Ericksen number, $\mathit{A}\mathrm{Er}$, with extensile activity ($A<0$) leading to parallel elongation and sufficient contractile activity ($A<0$) yielding perpendicular elongation [66], which is opposite to the aspect ratio’s depencend on the activity strength in the external problem.
Figures 4(c–e) show snapshots of the fluid stirring induced by extensile activity. Varying the activity and anchoring strengths primarily affects the speed of stirring, though contractile activity also reverses its direction. The maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, is plotted as function of the anchoring strength in Fig. 4(b), for both the LC external and internal to a cylinder. The speeds plateau for large anchoring strengths, $w\gtrsim 10$, and decrease according to $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ as $w\to 0$, for the external and internal problem, respectively. The extent to which this stirring flow results in mixing is an interesting but separate question [79]. Mixing is notoriously challenging in highly viscous flows and in confined geometries [80], but is likely promoted by a wide spatial distribution of active stress [35, 36].
Figure 4: (a) Deformation of a soft cylinder enclosing a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $w=10$, $M=50$, and $\kappa=5/3$. Color delineates the maximum shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the solid and the blue curves are director lines. (b) Plot of the maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, as a function of $w$ for the LC external and internal to a cylinder. The asymptotic behaviour as $w\to 0$, i.e. $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ (external) and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ (internal), are shown as black dashed lines. (c–e) Snapshots at times (c) $t=0$, (d) $t=100/|A|$, and (e) $t=200/|A|$ of colored fluid particles transported due to extensile activity inside the cylinder for $w=10$.
Conclusion:
We have shown that the configurations and dynamics of an active nematic can be affected dramatically by the mere presence of an immersed body, or boundary, for any non-trivial anchoring strength. The spontaneous flow that arises is characterized by plumes whose structure is anchoring strength dependent. The elastic, viscous, and active stresses may then deform the body, or external boundary, in a generally localized manner near (virtual) topological defects. Such flows can also produce a spontaneous stirring, and likely mixing, of the environment.
In the above theory, the deformation of the solid was assumed to be small enough that it does not affect the director field. In general, changes in the LC–solid interface could cause the director field to relax towards a different energy-minimizing configuration. Furthermore, although the flow does not directly influence the director field here, boundary deformations do depend on fluid flow, and so, this could provide an indirect influence if the interface deformation was taken into account: flows deform soft boundaries, thereby altering the LC. Even for a simple cylindrical domain, the deformation qualitatively depends on the activity strength (Fig. 3). Thus, since elongated colloids are known to reorient with the preferred axis of the surrounding LC and virtual defects seek out the corners of immersed bodies [70], significant changes to the LC could be expected here. The above work provides the first steps in modelling these complex fluid-structure interactions.
The localized deformations observed have also been reported in tactoids [78] and giant unilamellar vesicles [81]. Even for stiff bodies, the developed points could cause pinning of topological defects, which is important in setting the attraction angle of flexible colloid chains [82], for instance. Additionally, these deformations can encourage surface localized bundling of actin filaments [83, 84].
Looking to the horizon, biological cells are host to a multitude of active anisotropic networks. Can such fluids be triggered into motion solely by the passive anchoring conditions on an arriving organelle? This mechanism could offer an energy-efficient way to generate precisely timed mixing flows within the cellular environment.
Acknowledgements.
Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.

### v3

@@
 In an active nematic, the fluid components also generate a stress in the bulk fluid.
-By introducing either an immersed body or boundary, a fluid flow can be triggered due to anchoring boundary conditions.
-The fluid imposes elastic, viscous, and active stresses on such surfaces, which, if compliant, may result in a surface deformation.
-We study the deformations and stress in a linearly elastic body placed in an active nematic in two-dimensions.
+By introducing either an immersed body or boundary, a large scale flow can be triggered due to anchoring boundary conditions alone — a global pressure built by active stresses at equilibrium is instantly released everywhere.
+The fluid then imposes viscous, elastic, and active stresses on such surfaces which, if compliant, may result in a surface deformation.
+We study the deformations and stresses of a linearly elastic body placed in an active nematic in two dimensions.
 Using complex variables techniques, exact expressions for the fluid flow, director field, surface tractions, and body deformation are derived.
-Qualitative differences between elastic and active stress-driven deformations are identified, depending on an active Ericksen number, anchoring conditions, and body material properties, thereby suggesting a new method for measuring the active stresses in active anisotropic biological systems.
+Qualitative differences between elastic and active stress-driven deformations are identified, depending on an active Ericksen number, anchoring conditions, and body material properties, thereby suggesting a new method for measuring mechanical stresses in active anisotropic environments.
 Flow profiles, external confinement, and anchoring-induced stirring are also addressed.
+I Introduction
 The eukaryotic cell is host to a wide variety of deformable organelles, from the nucleus to mitochondria [1, 2, 3, 4, 5, 6].
-These organelles exist in a complex, potentially anisotropic, viscous environment [7, 8], and experience mechanical stresses due to actively driven cytoplasmic flows [9].
-The cell membrane itself is also deformable, and is commonly remodelled by internal processes.
-For instance, active anisotropic stresses are of great functional importance in cytoskeletal remodeling of the nucleus and chromatin configurations [10, 11], and in the metaphase spindle [12, 13, 14].
-On a larger length scale, active systems like bacterial swarms and biofilms also exhibit anisotropy, among other complex rheological properties [15, 16, 17, 18].
-Understanding how such bodies affect, and are affected by, this complex environment is of generic importance.
+These organelles reside in complex environments which can be anisotropic [7, 8], and experience mechanical stresses due to actively driven cytoplasmic flows [9, 10, 11, 12, 13, 14, 15].
+The cell membrane itself is also deformable, and its morphology is commonly influenced by internal processes.
+For instance, active anisotropic stresses are of functional importance in cytoskeletal remodeling of the nucleus and chromatin configurations [16, 17, 18, 19, 20], in the metaphase spindle [21, 22, 23], and in numerous other aspects of tissue morphogenesis [24, 25, 26].
+On a larger length scale, active systems like bacterial swarms and biofilms also exhibit anisotropy, among other complex rheological properties [27, 28, 29, 30, 31].
 Even when free of bodies, actively-driven anisotropic flows exhibit very rich and complex dynamics [32, 33, 34, 35, 36].
@@
 Confinement introduces an additional length scale, which can affect the emergent spatial structure [50, 51].
-Strong active nematic stresses can also change the shape of deformable confining surfaces [39, 40, 41, 42, 43, 44, 45, 46], and they play a prominent role in directing the motility of cells [47, 48] and active droplets [49, 50, 51, 52].
+Strong active nematic stresses can also change the shape of deformable confining surfaces [52, 53, 54, 55, 56, 57, 58, 59], and can play a prominent role in directing the motility of cells [60, 61] and active droplets [62, 63, 64, 65].
 When an active nematic contains an immersed body, or inclusion, additional elastic stresses are immediately introduced.
 In passive systems, these stresses result in intricate and beautiful textures [66, 67, 68], and can be used to measure cell material properties [69, 70].
-In the presence of activity, the combination of fluid elasticity and anisotropy redirects actively driven flows, leading to complex immersed body dynamics.
-Recent examples include an investigation of the biased rotation of gear-like bodies [58, 59], and transitions from fixed point to limit cycle to chaotic dynamics of a circular disk [60].
-How body deformability affects, and is affected, by such an active environment has not been well explored.
-In this letter, we present a model for an active anisotropic fluid with a soft internal or external boundary, confined to two-dimensions (2D).
-The model integrates Ericksen–Leslie theory [61, 62, 63] for active nematic liquid crystals (LCs), a common model for active anisotropic fluid [64, 23], and linear elasticity for deformable solids.
-Complex variable techniques yield novel analytical representations for the director field, velocity field, and boundary deformation.
-We demonstrate this for both an immersed and bounding cylinder, highlighting the insights that these analytical solutions provide.
-Active nematic LCs can be described by their locally-averaged molecular orientation, $\bm{n}(\bm{x},t)$, and fluid velocity $\bm{u}(\bm{x},t)$, with spatial position $\bm{x}$, time $t$, and $|\bm{n}|=1$.
-The director field, $\bm{n}$, evolves due to gradients in the velocity field, with strain rate tensor $\bm{\mathrm{E}}=(\nabla\bm{u}+\nabla\bm{u}^{T})/2$ and vorticity tensor $\bm{\Omega}=(\nabla\bm{u}-\nabla\bm{u}^{T})/2$, and a relaxation towards equilibrium under the action of the molecular field, $\bm{h}$.
-In the one-constant approximation, $\bm{h}=K\nabla^{2}\bm{n}$, for the single (Frank) elastic constant $K$.
-Altogether,
-$\partial_{t}\bm{n}+\bm{u}\cdot\nabla\bm{n}=\bm{n}\cdot\bm{\Omega}+\lambda\bm{n}\cdot\bm{\mathrm{E}}\cdot(\bm{\mathrm{I}}-\bm{n}\bm{n})+\tilde{\bm{h}}/\gamma,$ (1)
-with $\gamma$ rotational viscosity, $\lambda$ tumbling/reactive parameter, and $\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\bm{h}$ the transverse part of the molecular field, where $\bm{n}\bm{n}$ is a dyadic product [61, 62, 63].
-Assuming incompressibility, the fluid velocity satisfies the momentum and mass conservation equations
+In active systems, fluid elasticity and anisotropy can redirect the emergent flows, leading to complex body dynamics.
+Recent examples include an investigation of the biased rotation of gear-like bodies [71, 72], complex droplet trajectories [73, 74], and transitions from fixed point to limit cycle to chaotic dynamics of a circular disk [75].
+How body deformability affects and is affected by such an active environment remains largely unexplored.
+In this paper, we study an active nematic fluid based on Ericksen–Leslie theory [76, 77, 78, 79, 36] with a soft internal or external boundary in two dimensions.
+Complex variable techniques yield exact analytical representations for the director field, velocity field, and boundary deformation.
+Elastic deformations of bounding surfaces are known to relieve the elastic energy stored in the bulk LC phase [80].
+Here, we show that active stresses can compete and eventually dominate this balance, resulting in distinct geometric signals.
+In addition, we show that a spontaneous stirring flow with a unique plume structure may be induced with a change in surface anchoring boundary conditions alone, due to a global pressure release.
+II Active Nematic Flows
+Active nematic liquid crystals (LCs) can be described by their locally-averaged molecular orientation (i.e.
+the director field) $\bm{n}(\bm{x},t)$ and fluid velocity $\bm{u}(\bm{x},t)$, with spatial position $\bm{x}$, time $t$, and $|\bm{n}|=1$.
+The director field, $\bm{n}$, evolves due to gradients in the velocity field, with strain-rate tensor $E_{ij}=(\partial_{i}u_{j}+\partial_{j}u_{i})/2$ and vorticity tensor $\Omega_{ij}=(\partial_{i}u_{j}-\partial_{j}u_{i})/2$, and a relaxation towards equilibrium under the action of the molecular field, $\bm{H}=-\delta\mathcal{F}/\delta\bm{n}$, where $\mathcal{F}$ is the free energy density.
+In the one-constant approximation, $\mathcal{F}=K\lVert\nabla\bm{n}\rVert^{2}/2$ and so $\bm{H}=K\nabla^{2}\bm{n}$, for the single (Frank) elastic constant $K$ [77].
+Together,
+$(\partial_{t}+\bm{u}\cdot\nabla)\bm{n}=\bm{n}\cdot\bm{\Omega}+\lambda\bm{n}\cdot\bm{\mathrm{E}}\cdot(\bm{\mathrm{I}}-\bm{n}\bm{n})+\bm{h}/\gamma,$ (1)
+with $\gamma$ the rotational viscosity, $\lambda$ the tumbling/reactive parameter, and $\bm{h}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\bm{H}$ the transverse part of the molecular field, where $\bm{n}\bm{n}$ is a dyadic product [76, 77, 78, 80].
+Momentum and mass conservation, under the assumptions of incompressibility and zero Reynolds number flow [^1] The Reynolds number is defined as $\operatorname{Re}=\rho|\alpha|L^{2}/\mu^{2}$, with $\rho$ the fluid density, and other terms defined above., are given by
 $\nabla\cdot\left(\bm{\sigma}^{e}+\bm{\sigma}^{v}+\bm{\sigma}^{a}-p\bm{\mathrm{I}}\right)=\bm{0}\quad\text{and}\quad\nabla\cdot\bm{u}=0,$ (2a)
-for the pressure $p$ and elastic, viscous, and active stress tensors
-$\bm{\sigma}^{e}$ $=-K\nabla\bm{n}\cdot\nabla\bm{n}^{T}+\frac{1}{2}(1-\lambda)\bm{n}\bm{\tilde{h}}-\frac{1}{2}(1+\lambda)\bm{\tilde{h}}\bm{n},$ (3a)
+respectively, for $\nabla\cdot\bm{\sigma}\coloneqq\partial_{i}\sigma_{ij}$, pressure $p$, and elastic, viscous, and active stress tensors
+$\begin{split}\bm{\sigma}^{e}&=\frac{K}{2}\lVert\nabla\bm{n}\rVert^{2}\mathbf{I}-K\nabla\bm{n}\cdot\nabla\bm{n}^{T}\\ &\quad-\frac{1}{2}(\bm{n}\bm{h}-\bm{h}\bm{n})-\frac{\lambda}{2}(\bm{n}\bm{h}+\bm{h}\bm{n}),\end{split}$ (3a)
 […]
@@
 respectively.
-Here, $\mu$ is the solvent viscosity, $\mu_{1}$ and $\mu_{2}$ are anistropic viscosities, and $\alpha$ is the activity strength, which can describe both extensile ($\alpha<0$) and contractile ($\alpha>0$) activity [64, 65, 23].
-We non-dimensionalize the system by using an intrinsic length scale $L$ (e.g.
-the size of an immmersed body), velocity scale $U$ (to be defined), and stress scale $K/L^{2}$.
-The dimensionless system is then described by the dimensionless Ericksen number, activity strength, and viscosity ratios [66]:
+Here, $\mu$ is the solvent viscosity, $\mu_{1}$ and $\mu_{2}$ are anistropic viscosities, and $\alpha$ is the activity strength, which can describe both extensile ($\alpha<0$) and contractile ($\alpha>0$) activity [79, 81, 36] (see [82]).
+Note the convention taken for the divergence, which is important since $\bm{\sigma}^{e}$ is not symmetric.
+The system is made dimensionless by scaling upon an intrinsic length scale $L$ (e.g.
+the size of an immersed body), velocity scale $U$ (to be defined), and stress scale $K/L^{2}$.
+The system is then characterized by the dimensionless Ericksen number, activity strength, and viscosity ratios [82]:
 $\mathrm{Er}\coloneqq\frac{\mu UL}{K},\quad\mathit{A}\coloneqq\frac{\alpha L}{\mu U},\quad\gamma^{\prime}\coloneqq\frac{\gamma}{\mu},\quad\mu_{j}^{\prime}\coloneqq\frac{\mu_{j}}{\mu}.$ (4a)
-For small rotational Ericksen numbers, $\gamma^{\prime}\mathrm{Er}\ll 1$, the fluid flow does not affect the director field and the molecular field relaxes to equilibrium.
-The director evolution equation (1) takes the dimensionless form
-$\tilde{\bm{h}}=(\bm{\mathrm{I}}-\bm{n}\bm{n})\cdot\nabla^{2}\bm{n}=\mathcal{O}(\gamma^{\prime}\mathrm{Er}).$ (5)
-By further assuming small viscosity ratios, $\gamma^{\prime},\mu_{i}^{\prime}\ll 1$, the momentum equation (2a) takes the dimensionless form
-$\nabla^{2}\bm{u}-\nabla p=\frac{1}{\mathrm{Er}}\nabla\cdot\bigl{(}\nabla\bm{n}\cdot\nabla\bm{n}^{T}\bigr{)}-2\mathit{A}\nabla\cdot(\bm{n}\bm{n}),$ (6)
-up to $\mathcal{O}(\gamma^{\prime},\mu_{i}^{\prime})$. This describes an isotropic Stokes flow with active and elastic driving forces [66].
-In the examples presented in this letter, the fluid flow is induced by activity alone; thus, the velocity scale is set by the activity strength, $U=|\alpha|L/\mu$, and the assumption of a small Ericksen number is equivalent to assuming weak activity, $\gamma^{\prime}\mathrm{Er}=|\alpha|\gamma L^{2}/(\mu K)\ll 1$.
-Furthermore, in 2D, the assumption of a small viscosity ratio $\mu_{2}^{\prime}$ is unnecessary since it only corresponds to an anisotropy out-of-the-plane.
-Instead, $\mu_{2}$ can be absorbed into the isotropic viscosity by taking $\mu\mapsto\mu-\mu_{2}/2$.
+For small rotational Ericksen numbers, $\gamma^{\prime}\mathrm{Er}\ll 1$, the fluid flow does not affect the director field, and at leading order from (1) we have
+$\bm{h}=(\mathbf{I}-\bm{n}\bm{n})\cdot\nabla^{2}\bm{n}=\bm{0}.$ (5)
+Further, assuming small viscosity ratios, $\gamma^{\prime},\mu_{i}^{\prime}\ll 1$, Eq. (2a) yields
+$-\nabla p+\nabla^{2}\bm{u}+2\mathit{A}\nabla\cdot(\bm{n}\bm{n})=\bm{0}.$ (6)
+In the examples presented in this paper, the fluid flow is induced by activity alone; thus, the velocity scale is set by the activity strength, $U=|\alpha|L/\mu$, and the assumption of a small Ericksen number is equivalent to assuming weak activity, $\gamma^{\prime}\mathrm{Er}=|\alpha|\gamma L^{2}/(\mu K)\ll 1$.
+[^2] In 2D incompressible flows, a general $\mu_{2}$ can actually be absorbed into the isotropic viscosity by taking $\mu\mapsto\mu-\mu_{2}/2$.
 Figure 1: The activity-induced flow of a 2D nematic LC outside a cylinder for $\mathit{A}=-1$ (i.e.
@@
 Contractile activity with $\mathit{A}=1$ corresponds to flipping the flow direction.
-In two dimensions, the director field can be described by a director angle $\theta(x,y)\in[0,\pi)$ such that $\bm{n}=(\cos\theta,\sin\theta)$.
-Additionally, the conservation of mass (2b) implies that a streamfunction $\psi(x,y)$ can be introduced such that $\bm{u}=(\partial_{y}\psi,-\partial_{x}\psi)$.
-In this case, the equilibrium equation (5) reduces to Laplace’s equation, $\nabla^{2}\theta=0$, while the momentum equation (6) is equivalent to a Poisson’s equation for the scalar vorticity, $\omega=-\nabla^{2}\psi$.
-Here, we absorb the isotropic parts of the stress tensors into the pressure by taking $p\mapsto p+\mathit{A}-|\nabla\theta|^{2}/(2\mathrm{Er})$.
-At any boundaries, the fluid velocity is assumed to vanish, while the director relaxes to some preferential angle $\phi$ with some anchoring strength $W$ (i.e. Rapini–Papoular anchoring [67]):
-$\partial_{x}\psi=\partial_{y}\psi=0\hskip 6.00006pt\text{and}\hskip 6.00006pt\partial_{\nu}\theta=(w/2)\sin\left[2(\theta-\phi)\right],$ (7a)
-for the normal derivative pointing into the LC $\partial_{\nu}=\bm{\hat{\nu}}\cdot\nabla$ and the dimensionless anchoring strength $w\coloneqq WL/K.$
-Consider the complex position $z=x+\mathrm{i}y$ and its complex conjugate $\bar{z}=x-\mathrm{i}y$.
-Using the Wirtinger derivatives, $\partial_{x}=\partial_{z}+\partial_{\bar{z}}$ and $\partial_{y}=\mathrm{i}\partial_{z}-\mathrm{i}\partial_{\bar{z}}$, one can make a change of variables from $(x,y)$ to $(z,\bar{z})$ [68].
-In these complex coordinates, the system takes the form
+II.1 Two-dimensional active nematics
+In 2D, the director field can be described by an angle field $\theta(x,y)\in[0,\pi)$ such that $\bm{n}=(\cos\theta,\sin\theta)$, and a streamfunction $\psi(x,y)$ can be introduced such that $\bm{u}=(\partial_{y}\psi,-\partial_{x}\psi)$.
+The director field, governed by $\bm{h}=\bm{0}$, reduces to Laplace’s equation for $\theta$, while the momentum equation (6) is equivalent to Poisson’s equation for the scalar vorticity, $\omega=-\nabla^{2}\psi$:
+$\qquad\nabla^{2}\theta=0,$ (7a)
+$\nabla^{2}\omega=\mathit{A}\left[2\partial_{yx}\cos(2\theta)+(\partial_{yy}-\partial_{xx})\sin(2\theta)\right].$ (7b)
+Here, we absorb the isotropic parts of the stress tensors into the pressure by taking $p\mapsto p+\mathit{A}$. The leading-order, dimensionless, traceless stress tensors are denoted with hats:
+$\bm{\hat{\sigma}}^{e}$ $=\frac{1}{2}|\nabla\theta|^{2}\mathbf{I}-\nabla\theta\nabla\theta,$ (8a)
+$\bm{\hat{\sigma}}^{v}$ $=\nabla\bm{u}+\nabla\bm{u}^{T},$ (8b)
+$\bm{\hat{\sigma}}^{a}$ $=2\bm{n}\bm{n}-\mathbf{I},$ (8c)
+giving the dimensionless (bulk) stress tensor $\bm{\sigma}=\bm{\hat{\sigma}}^{e}+\mathrm{Er}\left(\bm{\hat{\sigma}}^{v}+A\bm{\hat{\sigma}}^{a}-p\mathbf{I}\right)$.
+At any boundaries, we assume the fluid velocity vanishes and the director field relaxes locally towards a preferred angle $\theta_{0}(s)$ due to surface anchoring of strength $W$ (i.e.
+Rapini–Papoular anchoring [83]):
+$\partial_{s}\psi=\partial_{\nu}\psi=0\quad\text{and}\quad\partial_{\nu}\theta=\frac{w}{2}\sin\left[2(\theta-\theta_{0})\right],$ (9a)
+for the anticlockwise arclength derivative $\partial_{s}=\bm{\hat{s}}\cdot\nabla$, normal derivative pointing into the LC $\partial_{\nu}=\bm{\hat{\nu}}\cdot\nabla$, and the dimensionless anchoring strength $w\coloneqq WL/K.$
+With $z=x+\mathrm{i}y$ and $\bar{z}=x-\mathrm{i}y$ its complex conjugate, the system may be written in complex variables as
 $\partial_{z\bar{z}}\theta=0\quad\text{and}\quad\partial_{\bar{z}}(p-\mathrm{i}\omega)=\mathit{A}\partial_{z}e^{2\mathrm{i}\theta},$ (10a)
@@
 for some locally-holomorphic functions $f$ and $g$ [84].
-(Throughout, we shall assume the LC does not contain any topological defects, and so, $f$ is non-zero and singularity free inside the LC.) Inserting the imaginary part of (9b) into $\partial_{z\bar{z}}\psi=-\omega/4$ and integrating yields
+We assume the LC does not contain any topological defects, so $f$ is non-zero and singularity free inside the LC, though this assumption can be relaxed [85, 86].
+Inserting the imaginary part of (11b) into $\partial_{z\bar{z}}\psi=-\omega/4$ and integrating yields
 $\psi=\frac{\mathit{A}}{4}\operatorname{Im}\Bigl{[}\bar{z}g(z)+h(z)+\overline{\int f(z)\mathop{}\!\mathrm{d}z}\,/f^{\prime}(z)\Bigr{]},$ (12)
 for another locally-holomorphic function $h$.
-The first two terms in (10) correspond to the well-known form of a biharmonic function with Goursat functions $h$ and $g$ [69].
-The last term is a particular solution that accounts for the active forcing.
-Analytical solutions for the director and velocity fields are thus obtained, provided that functions $f$, $g$, and $h$, which are locally holomorphic within the LC and satisfy the boundary conditions in (7), may be found.
+The first two terms in (12) correspond to the well-known form of a biharmonic function with Goursat functions $h$ and $g$ [87], while the last term accounts for the active forcing.
+Analytical solutions for the director and velocity fields are obtained, provided that functions $f$, $g$, and $h$, which are locally holomorphic within the LC and satisfy the boundary conditions in (9), may be found.
 The problem for $f$ derived here is equivalent to the potential problem explored in Refs.
-[70, 71] for a passive LC.
-It is important to note that the solutions (9)–(10) are not invariant under a conformal mapping since they contain a mixture of holomorphic (function of $z$) and antiholomorphic (function of $\bar{z}$) terms.
+[88, 89] for a passive LC, but $g$ and $h$ depend on activity.
+II.2 Anchoring-controlled flows
 As an example, consider a cylinder of radius $L$ of infinite extent immersed in a weakly-active nematic LC.
 In the far-field, we assume the LC is aligned horizontally in a direction orthogonal to the cylinder’s long axis and the fluid velocity vanishes.
-Additionally, the LC is assumed to be tangentially anchored to the cylinder with a finite strength, i.e.
-(7b) with $\phi=\arg(\mathrm{i}z)\mod\pi$ on $|z|=1$.
+The LC is assumed to be tangentially anchored to the body with finite strength, i.e.
+(9b) with $\theta_{0}=\arg(\mathrm{i}z)\mod\pi$ on $|z|=1$.
 Without the body, the system sits at an equilibrium, with $\bm{n}=\bm{\hat{x}}$ and $\bm{u}=\bm{0}$, and the active stress is absorbed by the pressure.
-First, we determine the LC director angle $\theta=-\arg f^{\prime}$.
+To find the activity-induced flow, we first determine the LC director angle $\theta=-\arg f^{\prime}$.
 Here, the potential $f(z)$ must be locally holomorphic in $1<|z|<\infty$ and satisfy the finite anchoring condition on $|z|=1$, and $f(z)\sim z$ as $|z|\to\infty$.
-The solution to this problem that minimizes the Frank elastic energy is, for any $w>0$,
+The solution to this problem is
 $\theta=-\arg f^{\prime}(z)=-\arg(1-\rho^{2}/z^{2}),$ (13)
-where $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$ is an effective radius [70]. The corresponding director lines (which lay parallel to $\bm{n}$) are plotted in Fig. 1(a) for $w=10$.
-Next, with (11), we determine the activity-induced fluid flow.
-Here, the Goursat functions $h(z)$ and $g(z)$ must be locally holomorphic in $1<|z|<\infty$ and yield a fluid velocity that vanishes on $|z|=1$ and as $|z|\to\infty$, ultimately yielding the streamfunction
-$\psi=\frac{\mathit{A}\rho^{2}}{8}\operatorname{Im}\biggl{[}\frac{(|z|^{2}-1)^{2}}{z^{2}(z^{2}-\rho^{2})}+\frac{1-|z|^{2}+2\log|z|}{(z^{2}-\rho^{2})/(2\rho^{2})}\biggr{]},$ (12)
-see [66].
-The corresponding velocity field is shown in Fig.
-1(b,c) for $w=1$ and $w=10$, with $\mathit{A}=-1$ (i.e.
-extensile activity).
-A prominent feature is a plume extending symmetrically to both the left and right of the cylinder.
-Far from the body, the velocity field behaves as
-$u+\mathrm{i}v\sim\bigl{[}\rho^{2}\cos(2\alpha)-\cos(4\alpha)\bigr{]}\mathit{A}\rho^{2}e^{\mathrm{i}\alpha}/(2|z|),$ (13)
-as $|z|\to\infty$, for $\alpha=\arg z$; the plume angle thus decreases from $\pi/4$ to 00 as the anchoring strength $w$ is increased from 00 to $\infty$, while the plume speed is largest at $w=8/3$.
-The velocity field does not appear as a dipole or stresslet in the far-field — the mere presence of the body initiates additional flow far away due to the distributed active stress, leading to bulk flow release.
+where $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$ is an effective radius [88].
+The corresponding director lines (which lay parallel to $\bm{n}$) are plotted in Fig.
+1(a) for $w=10$.
+Meanwhile, the Goursat functions $h(z)$ and $g(z)$ must be locally holomorphic in $1<|z|<\infty$ and yield a fluid velocity that vanishes on $|z|=1$ and as $|z|\to\infty$.
+This ultimately yields the streamfunction (see [82]):
+$\psi=\frac{\mathit{A}\rho^{2}}{8}\operatorname{Im}\biggl{[}\frac{(|z|^{2}-1)^{2}}{z^{2}(z^{2}-\rho^{2})}+\frac{1-|z|^{2}+2\log|z|}{(z^{2}-\rho^{2})/(2\rho^{2})}\biggr{]}.$ (14)
+The corresponding flow is shown in Fig. 1(b,c) for $w=1$ and $w=10$, with $\mathit{A}=-1$ (i.e. extensile activity).
+A prominent feature of the induced flow is a plume extending symmetrically to both the left and right of the cylinder (see Fig. 1(b,c)). Far from the body,
+$u+\mathrm{i}v=-2\mathrm{i}\partial_{\bar{z}}\psi\sim\bigl{[}\rho^{2}\cos(2\vartheta)-\cos(4\vartheta)\bigr{]}\frac{\mathit{A}\rho^{2}e^{\mathrm{i}\vartheta}}{2|z|},$ (15)
+as $|z|\to\infty$, for $\vartheta=\arg z$.
+This far-field flow behavior is critically dependent on the effective radius $\rho(w)$, and hence the anchoring strength $w$.
+The plume angle, thus, decreases from $\pi/4$ to 00 as the anchoring strength $w$ is increased from 00 to $\infty$, while the plume speed is largest at $w=8/3$.
+The velocity field does not appear as a dipole or stresslet in the far-field, but the sum of a dipole and a quadrupole — the mere presence of the body initiates additional flow far away due to the distributed active stress, leading to bulk flow release.
 This also demonstrates that a body with non-trivial anchoring conditions is sufficient to induce a large scale flow through its presence alone.
-Both the elasticity associated with passive LCs and the flow induced by activity generate stresses that act on any boundaries.
-Here, the passive LC introduces a surface traction $\bm{t}^{e}=\bm{\hat{\nu}}\cdot\hat{\bm{\sigma}}^{e}+\partial_{s}\bm{t}^{s}$, where $\hat{\bm{\sigma}}^{e}=(1/2)\lvert\nabla\theta\rvert^{2}\bm{\mathrm{I}}-\nabla\theta\nabla\theta$ is the (traceless) elastic stress tensor and $\bm{t}^{s}$ is a surface stress vector caused by the finite surface-anchoring (7b) [66].
-The activity generates an additional surface traction $\bm{t}^{a}=\mathrm{Er}\,\bm{\hat{\nu}}\cdot(\hat{\bm{\sigma}}^{v}+\mathit{A}\hat{\bm{\sigma}}^{a}-p\bm{\mathrm{I}})$ for the (traceless) viscous and active stress tensors $\hat{\bm{\sigma}}^{v}=\nabla\bm{u}+\nabla\bm{u}^{T}$ and $\hat{\bm{\sigma}}^{a}=2\bm{n}\bm{n}-\bm{\mathrm{I}}$, respectively.
+III Surface tractions and soft boundaries
+Stresses due to passive LC elasticity, actively-driven viscous flow, and direct activity all contribute to surface forcing.
+The LC introduces a surface traction $\bm{t}^{e}=\bm{\hat{\nu}}\cdot\bm{\hat{\sigma}}^{e}+\partial_{s}\bm{t}^{s}$, where $\bm{\hat{\sigma}}^{e}$ is the Ericksen elastic stress tensor given in (8) and
+$\bm{t}^{s}=\frac{w}{2}\sin^{2}(\theta-\theta_{0})\bm{\hat{s}}+\frac{w}{2}\sin[2(\theta-\theta_{0})]\bm{\hat{\nu}},$ (16)
+is a surface stress vector associated with finite surface-anchoring [80].
+Activity, meanwhile, produces a surface traction $\bm{t}^{a}=\mathrm{Er}\,\bm{\hat{\nu}}\cdot(\bm{\hat{\sigma}}^{v}+\mathit{A}\bm{\hat{\sigma}}^{a}-p\bm{\mathrm{I}})$ for the (traceless) viscous and active stress tensors, $\bm{\hat{\sigma}}^{v}$ and $\bm{\hat{\sigma}}^{a}$ given in (8).
 If a boundary is soft, it will generally be deformed by the surface tractions above.
-We assume the boundary is part of an isotropic solid that deforms according to plane stress/strain in the framework of linear elasticity, with shear modulus $\mu_{s}$ and Poisson’s ratio $\nu$ [72, 73].
-In this framework, the deformed (Eulerian) frame and the reference (Lagrangian) frame are mathematically equivalent.
-At equilbrium, the elastic solid can be described by an Airy-stress function, $\mathcal{A}(x,y)$, which satisfies the biharmonic equation, $\nabla^{4}\mathcal{A}=0$.
+As a first step towards understanding more complex materials, we assume the boundary is part of an isotropic solid that deforms according to plane stress/strain in the framework of linear elasticity, with shear modulus $\mu_{s}$ and Poisson’s ratio $\nu$ [90, 91].
+At equilibrium, the elastic solid can be described by an Airy-stress function, $\mathcal{A}(x,y)$, which satisfies the biharmonic equation, $\nabla^{4}\mathcal{A}=0$.
 We can, thus, write
@@
 for two (new) Goursat functions, $G$ and $H$, that are locally holomorphic within the solid (but not necessarily the LC), analogous to $g$ and $h$ in (12) [90].
-These Goursat functions yield the stress tensor within the solid, $\bm{\Sigma}$, via the Kolosov–Muskhelishvili formulae [66].
-Additionally, the solid displacement, $(U,V)$, can be expressed as
+These Goursat functions yield the symmetric stress tensor within the solid, $\Sigma_{ij}$, via the Kolosov–Muskhelishvili formulae,
+$\Sigma_{11}-\Sigma_{22}-2\mathrm{i}\Sigma_{12}=2\mathrm{i}\bar{z}G^{\prime\prime}(z)+2\mathrm{i}H^{\prime\prime}(z)$ (18a)
+$\text{and}\quad\Sigma_{11}+\Sigma_{22}=4\operatorname{Im}G^{\prime}(z),$ (18b)
+and the solid displacement, $(U,V)$, can be expressed as
 $2M(U-\mathrm{i}V)=\mathrm{i}\kappa\overline{G(z)}+\mathrm{i}\bar{z}G^{\prime}(z)+\mathrm{i}H^{\prime}(z),$ (19)
-up to an arbitrary body displacement and rotation, for the dimensionless elastic modulus $\mathit{M}\coloneqq\mu_{s}L^{2}/K$ and $\kappa=(3-4\nu)$ or $(3-\nu)/(1+\nu)$ for plane strain or plane stress, respectively.
-The LC–solid interface must also be at equilibrium, and so, the stresses acting on it must balance: $-\bm{\hat{\nu}}\cdot\bm{\Sigma}=\bm{t}^{e}+\bm{t}^{a}$.
-Provided we can find functions $G$ and $H$ that are locally holomorphic within the solid and satisfy this stress balance, we will have obtained analytic solutions for the solid’s deformation.
-However, yet again, the above solution is not invariant under a conformal map, and analytical progress can only be made in specific cases.
-We return to the example of a unit cylinder subject to (finite) tangential anchoring.
-Here, the director angle (11) and streamfunction (12) yield analytic expressions for the elastic traction, $t^{e}_{x}+\mathrm{i}t_{y}^{e}$, and activity-induced traction, $t^{a}_{x}+\mathrm{i}t_{y}^{a}=\mathit{A}\mathrm{Er}(1/z-\rho^{2}z)$, acting on the cylinder.
+up to an arbitrary body displacement and rotation, for the dimensionless elastic modulus $\mathit{M}\coloneqq\mu_{s}L^{2}/K$ and $\kappa=(3-4\nu)$ or $(3-\nu)/(1+\nu)$, when assuming plane strain or plane stress, respectively.
+Stress balance at the LC-solid interface, $\bm{\hat{\nu}}\cdot\bm{\Sigma}+\bm{t}^{e}+\bm{t}^{a}=\bm{0}$, yields the boundary conditions for $G$ and $H$:
+$\partial_{s}\left[\bar{z}G^{\prime}(z)-\overline{G(z)}+H^{\prime}(z)\right]=(t^{e}_{x}+t^{a}_{x})-\mathrm{i}(t^{a}_{y}+t^{e}_{y}).$ (20)
+III.1 Activity-controlled deformations
+Returning to the example of a unit cylinder with tangential anchoring, the director angle (13) and streamfunction (14) yield analytic expressions for the elastic traction, $t^{e}_{x}+\mathrm{i}t_{y}^{e}$, and activity-induced traction, $t^{a}_{x}+\mathrm{i}t_{y}^{a}=\mathit{A}\mathrm{Er}(1/z-\rho^{2}z)$.
 The cylinder is now assumed to be deformable, and so, we seek the Airy stress (17).
-The Goursat functions $G(z)$ and $H(z)$ must be locally holomorphic in $|z|<1$ and satisfy the stress balance condition on $|z|=1$.
-See [66] for the solution to this problem.
-For passive LCs ($\mathit{A}=0$), the deformation acts to reduce the elastic energy stored in the LC by flattening the cylinder.
-Figure  2 shows the LC configurations and body deformations for three different anchoring strengths and $\kappa=5/3$ (i.e.
+Here, the Goursat functions $G(z)$ and $H(z)$ must be locally holomorphic in $|z|<1$ and satisfy (20) on $|z|=1$; the solution is provided in [82].
+Without activity ($\mathit{A}=0$), boundary deformations help to reduce the elastic energy stored in the bulk LC [80].
+Figure 2 shows the LC configurations and body deformations for three different anchoring strengths and $\kappa=5/3$ (i.e.
 an incompressible cylinder, $\nu=1/2$, under plane stress).
 For large anchoring strengths, $w\gg 1$, most of the LC energy is stored local to the left and right poles of the cylinder, with a singular energy emerging in the case of strong (infinite) anchoring ($w=\infty$) due to the appearance of two $-1$ Boojum topological defects at $z=\pm 1$ [92, 88].
-This localization of energy within the LC results in large localized surface tractions: $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(w^{2})$ for $z\pm 1=\mathcal{O}(1/w)$, and $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(1)$ otherwise, as $w\to\infty$.
-This, in turn, leads to large localized deformations of the cylinder, see Fig.
-2.
-In particular, we find that
+This localized energy within the LC results in large localized surface tractions as $w\to\infty$: $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(w^{2})$ when $z\pm 1=\mathcal{O}(1/w)$, and $t^{e}_{x}+\mathrm{i}t_{y}^{e}=\mathcal{O}(1)$ otherwise.
+This, in turn, leads to large localized deformations; in particular, we find that
 $U+\mathrm{i}V\sim\frac{3w}{8M}\bigl{[}\mathcal{U}(1-z)-\mathcal{U}(1+z)\bigr{]},$ (21)
 as $w\to\infty$, where
-$\mathcal{U}(Z/w)=\kappa/(1+Z)+1/(1+\bar{Z})+(Z+\bar{Z})/(1+\bar{Z})^{2},$ (17)
-corresponds to a deformation that is only apparent local to the left and right poles, i.e.
-when $z\pm 1=\mathcal{O}(1/w)$.
-As the anchoring strength decreases, stress is no longer concentrated at the poles and the maximum shear stress $\tau(x,y)=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$ (a common indicator of mechanical failure [73]) is greatest at (±$(\pm x_{\max},0)$, where $x_{\max}$ decreases from $1$ for $w=\infty$, to 00 for $w\leq 24/35$, see Fig.
+$\mathcal{U}(Z/w)=\frac{\kappa}{1+Z}+\frac{1}{1+\bar{Z}}+\frac{Z+\bar{Z}}{(1+\bar{Z})^{2}},$ (22)
+corresponds to a deformation that is only apparent when $z\pm 1=\mathcal{O}(1/w)$.
+As the anchoring strength decreases, stress is no longer concentrated at the poles and the maximum shear stress $\tau(x,y)=\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$ (a common indicator of mechanical failure [91]) is greatest at (±$(\pm x_{\max},0)$, where $x_{\max}$ decreases from $1$ when $w=\infty$, to 00 when $w\leq 24/35$, as shown in Fig.
 2.
-Figure 2: (top) Deformation of soft cylinders immersed in passive nematic LCs with $\mathit{A}\mathrm{Er}=0$, $\mathit{M}/w=5$, $\kappa=5/3$, and dimensionless anchoring strengths ① $w=0.1$, ② $w=1$, and ③ $w=10$.
-Color indicates the maximum shear stress, $\tau=\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines.
-(bottom) The positions, $(x_{\max},0)$, and value, $\tau_{\max}$, of the largest maximum shear stress plotted as a function of $w$.
-For active LCs ($A\neq 0$), the activity leads to an additional deformation:
-$U+\mathrm{i}V=U^{e}+\mathrm{i}V^{e}+\mathit{A}\mathrm{Er}\bigl{[}(1-\kappa)\rho^{2}z+2\bar{z}\bigr{]}/(4\mathit{M}),$ (18)
+Figure 2: (top) Deformation of a soft cylinder immersed in a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $\mathit{M}/w=5$, $\kappa=5/3$, and dimensionless anchoring strengths ① $w=0.1$, ② $w=1$, and ③ $w=10$.
+Color indicates the maximum shear stress, $\tau=\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the cylinder and the blue curves are director lines.
+(bottom) The positions, $(x_{\max},0)$, and value, $\tau_{\max}=\max_{\bm{x}}\tau$, of the largest maximum shear stress plotted as a function of $w$.
+In an active nematic ($A\neq 0$), the activity leads to an additional deformation:
+$U+\mathrm{i}V=U^{e}+\mathrm{i}V^{e}+\frac{\mathit{A}\mathrm{Er}}{4\mathit{M}}\bigl{[}(1-\kappa)\rho^{2}z+2\bar{z}\bigr{]},$ (23)
 where $(U^{e},V^{e})$ is the deformation of a cylinder immersed in a passive LC.
@@
 In particular, the aspect ratio, $\mathcal{A}_{R}$, of the deformed cylinder is given by
-$\mathcal{A}_{R}\sim 1+U\bigr{|}_{z=1}-V\bigr{|}_{z=\mathrm{i}}=1+(\mathit{A}\mathrm{Er}-\mathcal{W})/\mathit{M},$ (19)
+$\mathcal{A}_{R}\sim 1+U\bigr{|}_{z=1}-V\bigr{|}_{z=\mathrm{i}}=1+\frac{\mathit{A}\mathrm{Er}-\mathcal{W}}{\mathit{M}},$ (24)
 where
@@
 Therefore, contractile activity ($\mathit{A}>0$), in this setting, always leads to parallel elongation, while extensile activity ($\mathit{A}<0$) yields perpendicular elongation, as long as the activity is sufficiently strong.
-Figure 3: (top) Deformation of soft cylinders immersed in active nematic LCs with $w=5$, $\mathit{M}=50$, $\kappa=5/3$, and ④ $\mathit{A}\mathrm{Er}=-8<\mathcal{W}$, ⑤ $\mathit{A}\mathrm{Er}=\mathcal{W}\approx-3.55$, and ⑥ $\mathit{A}\mathrm{Er}=4>\mathcal{W}$.
-Color indicates the maximal shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the cylinder and the blue curves are director lines.
+Figure 3: (top) Deformation of a soft cylinder immersed in an active nematic LC with $w=5$, $\mathit{M}=50$, $\kappa=5/3$, and ④ $\mathit{A}\mathrm{Er}=-8<\mathcal{W}$ (extensile), ⑤ $\mathit{A}\mathrm{Er}=\mathcal{W}\approx-3.55$ (extensile), and ⑥ $\mathit{A}\mathrm{Er}=4>\mathcal{W}$ (contractile).
+Color indicates the maximal shear stress, $\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the cylinder and the blue curves are director lines.
 (bottom) The critical active Ericksen number, $\mathit{A}\mathrm{Er}=\mathcal{W}$, at which the deformed cylinder has aspect ratio one, plotted as a function of $w$ for $\kappa=5/3$.
@@
 2.
-The fluid flow triggered by the presence of a boundary may also be advantageous in biological cells, settings in which a disperse distribution of molecules, such as nutrients, chemicals, or oxygen, benefit the system’s function [75, 76, 77].
-Consider an active nematic LC bounded inside, and tangentially anchored to, a cylinder with anchoring strength $w$.
-Here, the director angle of the LC can be written as
-$\theta=\arg(1-\rho^{2}z^{2}),$ (21)
-for the effective radius introduced in the external problem, $\rho(w)\coloneqq({\sqrt{1+4/w^{2}}-2/w})^{1/2}$, while the activity-induced flow can be expressed in terms of a streamfunction
-$\begin{split}\psi=\frac{\mathit{A}}{8}(|z|^{2}-1)\operatorname{Im}\Big{[}&\frac{1-\rho z}{z^{2}}\log(1-\rho z)\\ +&\frac{1+\rho z}{z^{2}}\log(1+\rho z)\Big{]},\end{split}$ (22)
-see [66].
-As before, the elasticity, activity, and viscosity of the LC will induce surface stresses on the boundary, which, if compliant, will deform.
+III.2 Bounded domains and spontaneous stirring
+The fluid flow triggered by the presence of a boundary may also be advantageous in interior domains like biological cells — settings in which a distribution of molecules, such as nutrients, chemicals, or oxygen, benefit the system’s function [93, 94, 95, 96].
+Consider an active nematic LC bounded inside and (weakly) tangentially anchored to a cylinder with anchoring strength $w$.
+Here, the director angle of the LC can be written as $\theta=\arg(1-\rho^{2}z^{2})$, for the effective radius introduced in the external problem, $\rho(w)=({\sqrt{1+4/w^{2}}-2/w})^{1/2}$, while the flow can be expressed in terms of a streamfunction (see [82]):
+$\begin{split}\psi=\frac{\mathit{A}}{8}(|z|^{2}-1)\operatorname{Im}\Big{[}\frac{1-\rho z}{z^{2}}&\log(1-\rho z)\\ +&\frac{1+\rho z}{z^{2}}\log(1+\rho z)\Big{]}.\end{split}$ (26)
 Figure 4(a) shows the director lines inside a cylinder that has deformed according to linear elasticity [82].
 The final deformed shape resembles a tactoid [97], with a localized deformation at the left and right poles occurring for large anchoring strengths due to the appearance of two $+1$ aster defects when $w=\infty$.
-The deformed shape is dependent on the active Ericksen number, $\mathit{A}\mathrm{Er}$, with extensile activity ($A<0$) leading to parallel elongation and sufficient contractile activity ($A<0$) yielding perpendicular elongation [66], which is opposite to the aspect ratio’s depencend on the activity strength in the external problem.
-Figures 4(c–e) show snapshots of the fluid stirring induced by extensile activity.
-Varying the activity and anchoring strengths primarily affects the speed of stirring, though contractile activity also reverses its direction.
-The maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, is plotted as function of the anchoring strength in Fig.
-4(b), for both the LC external and internal to a cylinder.
-The speeds plateau for large anchoring strengths, $w\gtrsim 10$, and decrease according to $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ as $w\to 0$, for the external and internal problem, respectively.
-The extent to which this stirring flow results in mixing is an interesting but separate question [79].
+The deformed shape is dependent on the active Ericksen number, $\mathit{A}\mathrm{Er}$, with extensile activity ($A<0$) leading to parallel elongation and sufficient contractile activity ($A<0$) yielding perpendicular elongation [82] — opposite to the dependence on the activity strength in the external problem.
+Figures 4(c–e) show stirring induced by extensile activity.
+Varying the activity and anchoring strength affects the rate of stirring.
+The maximum fluid speed, $|\bm{u}|_{\max}$, is plotted as function of anchoring strength in Fig.
+4(b), for both the external and internal problems.
+The speeds plateau for large anchoring strengths, and decrease according to $|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ as $w\to 0$, for the external and internal problem, respectively.
+The extent to which this stirring flow is mixing is an interesting but separate question [98].
 Mixing is notoriously challenging in highly viscous flows and in confined geometries [99], but is likely promoted by a wide spatial distribution of active stress [48, 49].
 Figure 4: (a) Deformation of a soft cylinder enclosing a passive nematic LC with $\mathit{A}\mathrm{Er}=0$, $w=10$, $M=50$, and $\kappa=5/3$.
-Color delineates the maximum shear stress, $\sqrt{(\Sigma_{xx}-\Sigma_{yy})^{2}/4+\Sigma_{xy}^{2}}$, within the solid and the blue curves are director lines.
+Color delineates the maximum shear stress, $\sqrt{(\Sigma_{11}-\Sigma_{22})^{2}/4+\Sigma_{12}^{2}}$, within the solid and the blue curves are director lines.
 (b) Plot of the maximum fluid speed, $|\bm{u}|_{\max}=\max_{\bm{x}}\sqrt{u^{2}+v^{2}}$, as a function of $w$ for the LC external and internal to a cylinder.
-The asymptotic behaviour as $w\to 0$, i.e.
-$|\bm{u}|_{\max}\sim 2|\mathit{A}|w/(25\sqrt{5})$ (external) and $|\bm{u}|_{\max}\sim|\mathit{A}|w^{2}/(600\sqrt{5})$ (internal), are shown as black dashed lines.
+The asymptotic behaviours as $w\to 0$ are shown as black dashed lines.
 (c–e) Snapshots at times (c) $t=0$, (d) $t=100/|A|$, and (e) $t=200/|A|$ of colored fluid particles transported due to extensile activity inside the cylinder for $w=10$.
-Conclusion:
-We have shown that the configurations and dynamics of an active nematic can be affected dramatically by the mere presence of an immersed body, or boundary, for any non-trivial anchoring strength.
-The spontaneous flow that arises is characterized by plumes whose structure is anchoring strength dependent.
-The elastic, viscous, and active stresses may then deform the body, or external boundary, in a generally localized manner near (virtual) topological defects.
+IV Conclusions and discussions
+We have shown that the configurations and dynamics of an active nematic can be affected dramatically by the mere presence of a boundary, for any non-trivial anchoring strength.
+It suggests that energy bound in a uniform active nematic can be released as a large scale flow due to a localized inclusion alone.
+The spontaneous flow that arises is characterized by unique plumes whose structure is anchoring strength dependent.
+The body may then deform in response to this anisotropic, viscoelastic flow, in a generally localized manner near (virtual) topological defects.
 Such flows can also produce a spontaneous stirring, and likely mixing, of the environment.
-In the above theory, the deformation of the solid was assumed to be small enough that it does not affect the director field.
-In general, changes in the LC–solid interface could cause the director field to relax towards a different energy-minimizing configuration.
-Furthermore, although the flow does not directly influence the director field here, boundary deformations do depend on fluid flow, and so, this could provide an indirect influence if the interface deformation was taken into account: flows deform soft boundaries, thereby altering the LC.
-Even for a simple cylindrical domain, the deformation qualitatively depends on the activity strength (Fig.
-3).
-Thus, since elongated colloids are known to reorient with the preferred axis of the surrounding LC and virtual defects seek out the corners of immersed bodies [70], significant changes to the LC could be expected here.
-The above work provides the first steps in modelling these complex fluid-structure interactions.
-The localized deformations observed have also been reported in tactoids [78] and giant unilamellar vesicles [81].
-Even for stiff bodies, the developed points could cause pinning of topological defects, which is important in setting the attraction angle of flexible colloid chains [82], for instance.
-Additionally, these deformations can encourage surface localized bundling of actin filaments [83, 84].
+The associated symmetry breaking can lead to cell or droplet division and motility [100, 101, 102].
 Looking to the horizon, biological cells are host to a multitude of active anisotropic networks.
@@
 This mechanism could offer an energy-efficient way to generate precisely timed mixing flows within the cellular environment.
+New experiments on the shapes and dynamics of deformable inclusions in active nematics is needed.
+Testable hypotheses include the differential direction of elongation, which may be tuned by varying either the activity strength (e.g.
+by changing the density of molecular motors  [23]) or by changing the body stiffness.
+With sufficient perpendicular elongation due to strong extensile activity, as in the top left panel of Fig.
+3, body rotation may also ensue to reduce the bulk elastic energy.
+We thus predict that a diverse zoology of body dynamics will emerge with increasing extensile activity in particular.
+[…]
+[…]
+Future directions will require different machinery.
+Large deformations will more completely couple the LC configuration and flows, and a finite rotational viscosity will introduce an additional timescale [77, 109].
+The analytical limits provided herein may still be of use for steering investigations of these more highly nonlinear regimes.
 Acknowledgements.
+We acknowledge helpful discussions with Nicholas L.
+Abbott and Ido Lavi.
 Support for this research was provided by the Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation.